OpenStudy (anonymous):

Find the Mistake: provided below is a solution to the equation 3 − 4x = 10x + 10. Choose the statement that is true about the given solution. Line # 3 − 4x = 10x + 10 1 3 = 14x + 10 2 −7 = 14x 3 −2 = x

OpenStudy (anonymous):

Please explain it to me (:

OpenStudy (cwrw238):

staatemnt 3 is incorrect -7 = 14x divide both sides by 14 -7/14 = x -1/2 = x

OpenStudy (anonymous):

Thank you!